what is the perimeter of the golden rectangle that can be constructed in a square whose side is 6 meter?
Answer by stanbon(75887)

what is the perimeter of the golden rectangle that can be constructed in a square whose side is 6 meter?
------------
If the longer side is 6 meters the shorter side is 6/1.618 = 3.7083
---
Then perimeter = 2(6+3.7083) = 19.4166 meters
=============================================